# IEPile: Unearthing Large-Scale Schema-Based Information Extraction Corpus

### v2

@@ B.3 Data Statistics
Based on the aforementioned methodologies, we construct a high-quality information extraction instruction dataset known as IEPile. This dataset contains approximately two million instances, each comprising two fields: “instruction” and “output”, which are formatted for direct use in the supervised fine-tuning of models. In terms of storage, IEPile occupies around 3GB of disk space, encompassing roughly 0.32B tokens (utilizing the BaiChuan2 tokenizer).
@@ C.1 Experimental Settings
For evaluating zero-shot generalization capabilities, we select a range of strong models for comparative analysis:
@@
Zero-shot Dataset
To ensure the validity of the zero-shot evaluation and prevent result bias due to data similarity, we select datasets primarily derived from news and biomedical fields as our training sets. This selection is intended to train the model’s capability for instruction following and schema-based extraction. For the evaluation data, we adopt the 13 cross-domain datasets recommended in IEINSTRUCTIONS and YAYI-UIE, which include: for Named Entity Recognition (NER) tasks, we use the CrossNER  (Liu et al., 2021), Weibo NER  (Peng and Dredze, 2015), and Boson [^4] https://github.com/InsaneLife/; in Relation Extraction (RE) tasks, we involve FewRel  (Han et al., 2018), Wiki-ZSL  (Chen and Li, 2021), COAE2016555https://github.com/Sewens/COAE2016, IPRE  (Wang et al., 2019), and SKE2020666https://aistudio.baidu.com/datasetdetail/177191; and for Event Extraction (EE), we include RAMS  (Ebner et al., 2020), WikiEvents  (Li et al., 2021), CrudeOilNews  (Lee et al., 2022b), FewFC  (Zhou et al., 2021), and CCF law  [^7] https://aistudio.baidu.com/projectdetail/4201483. These datasets cover a wide range of fields including literature, music, law, and oil news. It is noteworthy that these evaluation data sets are not used during the training, ensuring that our evaluation accurately reflects the model’s generalization and adaptation capabilities for unseen domains and unseen schema data in zero-shot information extraction.
Hyperparameter Value
Number of Epochs 5
Learning Rate 5e-5
Batch Size 20
Accumulate 4
Optimizer Adam
Lora_r 64
Lora_alpha 64
Lora_dropout 0.05
Table 4: Training Hyperparameters
@@
C.2 Zero-shot performance on Event Extraction
@@
As illustrated in Table  5, the model trained with IEPile exhibits outstanding performance in zero-shot event extraction (EE) tasks, surpassing other competing models. Notably, in the Chinese EE task, the LLaMA2-IEPile model’s performance is slightly inferior to YAYI-UIE’s, revealing LLaMA2’s limitations in processing Chinese data. However, in the English EE task, LLaMA2-IEPile’s performance is significantly superior to that of similar models. This contrast highlights the potential influence of language type on model performance.
C.3 Hyper-parameter
@@
In our research, we select two pre-trained models, Baichuan2-13B-Chat and LLaMA2-13B-Chat, as the base models for our study. Specifically, we employ the LoRA  (Hu et al., 2022) technique and utilize 8 NVIDIA A800 GPUs to perform instruction tuning on our IEPile dataset. The training period for each model spans approximately three days. Detailed configurations of the hyperparameters during the fine-tuning process are presented in Table 4.
Dataset InstructUIE YAYI-UIE Baichuan2-IEPile LLaMA2-IEPile
ACE2005 86.66 81.78 81.86 81.14
AnatEM 90.89 76.54 87.21 86.90
BC2GM 85.16 82.05 80.73 83.07
BC4CHEMD 90.30 88.46 90.45 90.07
BC5CDR 89.59 83.67 88.07 88.01
CoNLL2003 92.94 96.77 92.49 92.98
FabNER 76.20 72.63 77.07 76.33
FindVehicle 89.47 98.47 98.49 97.91
GENIA-Ent 74.71 75.21 76.66 77.32
HarveyNER 88.79 69.57 67.70 62.64
MIT Movie 89.01 70.14 88.23 89.54
MIT Restaurant 82.55 79.38 79.85 81.30
MultiNERD 92.32 88.42 94.60 94.24
NCBI-Disease 90.23 87.29 85.26 87.59
Ontonotes 90.19 87.04 87.55 90.34
Avg 87.27 82.49 85.08 85.29
MSRA - 95.57 87.99 86.32
Resume NER - - 93.92 92.86
CLUE NER - - 80.19 76.57
@@
Dataset InstructUIE YAYI-UIE Baichuan2-IEPile LLaMA2-IEPile
ADE Corpus 82.31 84.14 83.73 85.87
CoNLL2004 78.48 79.73 72.87 73.71
GIDS 81.98 72.36 74.71 74.13
KBP37 36.14 59.35 65.09 61.49
NYT 90.47 89.97 93.0 92.22
NYT11-HRL 56.06 57.53 53.19 54.86
SciERC 45.15 40.94 43.53 44.58
Semeval-RE 73.23 61.02 58.47 57.61
Avg 67.98 68.13 68.07 68.06
CMeIE - - 49.16 47.40
DuIE2.0 - 81.19 75.61 74.34
@@
Dataset InstructUIE YAYI-UIE Baichuan2-IEPile LLaMA-IEPile
Trigger ACE2005 77.13 65.00 72.46 70.63
CASIE 67.80 63.00 60.07 61.27
PHEE 70.14 63.00 66.22 68.52
Avg 71.69 63.67 66.25 66.81
DuEE1.0 - 85.00 86.73 84.01
DuEE-fin - 82.50 83.54 79.00
Argument ACE2005 72.94 62.71 63.90 62.69
CASIE 63.53 64.23 56.07 56.78
PHEE 62.91 77.19 70.85 71.33
Avg 66.46 68.04 63.60 63.61
DuEE1.0 - 79.08 75.63 73.79
DuEE-fin - 70.02 79.34 73.08
@@
C.4 Supervision Results
Tables  6,  7, and  8 present our experimental results under a supervised learning setting on the training dataset. Specifically, it can be observed that after training on the IEPile, the model excels in Named Entity Recognition (NER), Relation Extraction (RE), and Event Detection (ED), ranking second across these tasks. The model’s performance is only slightly behind other reference models in the Event Argument Extraction. Additionally, we record the model’s performance in Chinese NER, RE, and EE tasks, where it demonstrates robust results. Notably, unlike fully parameter-fine-tuned models such as InstructUIE and YAYI-UIE, we employ the Low-Rank Adaptation (LoRA) technique for model optimization, which involves partial retraining. This difference in fine-tuning approach may have a certain impact on the results of supervision learning. In a comprehensive assessment, the IEPile-trained model showcases performance on par with other models in instruction-based information extraction (IE) tasks and significantly improves performance in zero-shot IE tasks compared to other models. This indicates the significant application prospects and potential of IEPile in the current field of IE.
C.5 Impact of Potential Dataset Bias on Model Performance and Generalization
During the IEPile project, we identify that potential biases introduced by the datasets used can affect the model’s performance and generalization capability. Firstly, biases in the definition of schemas within the datasets have a negative impact on model performance  (Huang et al., 2024). In the early stages of training, we observe instability in results due to mutual interference among multiple datasets that contain the same schemas but with differing definitions (such as wikiann, wikineural, polyglot-NER, and CoNLL2003). Consequently, in the later stages, we remove all such datasets except for CoNLL2003. Secondly, the model demonstrates good generalization when dealing with datasets having schemas similar to those in the training set. For instance, the model proficiently handles People Daily and ACE2004 datasets, which share similar schemas to the MASR and ACE2005 NER datasets in the training set, despite not being included in the training corpus. Lastly, the use of common, coarse-grained labels (such as “person” and “organization”) within the IEPile project lead the model, after training, to favor these coarse categories over fine-grained ones (such as “scientist” and “company”) when predicting instructions that included both levels of granularity.
@@ C.5 Impact of Potential Dataset Bias on Model Performance and Generalization
Table 9: Statistical data of Named Entity Recognition (NER) datasets, with an asterisk ($\dagger$) indicating a zero-shot evaluation set not included in the instruction tuning. CrossNER (Liu et al., 2021) is divided into five subsets for our statistical analysis. For the zero-shot evaluation sets, we only report their test set.
@@
COAE2016$\dagger$ 5 General 9 - - 971
@@
SKE2020$\dagger$ 6 News 49 - - 3601
Table 10: Statistical data of Relation Extraction (RE) datasets, with an asterisk ($\dagger$) indicating a zero-shot evaluation set not included in the instruction tuning. The test sets for CMeIE and DuIE2.0 are not open-sourced, thus we use the validation sets as our evaluation data. For the FewRel and Wiki-ZSL datasets, we follow  Chia et al. (2022).
@@
CCF law $\dagger$7 Law 9(39) - - 971
Table 11: Statistical data of Event Extraction (EE) datasets, with an asterisk ($\dagger$) indicating a zero-shot evaluation set not included in the instruction tuning. The test sets for DuEE1.0 and DuEE-Fin are not open-sourced, thus we use the validation sets as our evaluation data.

### v3

@@ B.3 Data Statistics
Based on the aforementioned methodologies, we construct a high-quality information extraction instruction dataset known as IEPile. This dataset contains approximately two million instances and approximately 0.32B tokens. Each instance of IEPile comprises two fields: “instruction” and “output”, which are formatted for direct use in the instruction tuning.
@@ C.1 Experimental Settings
To assess the zero-shot generalization capabilities, we select a range of strong models for comparative analysis:
@@
  • GPT-4  (OpenAI, 2023): Known as the most powerful closed-source chat model to date.
  • LLaMA3  [^4] https://ai.meta.com/blog/meta-llama-3/.: The latest release in the LLaMA model series, achieving significant improvements across various benchmarks.
@@
C.2 OneKE
We leverage IEPile, InstructIE (Gui et al., 2023), CMRC (Cui et al., 2019), along with certain proprietary business information extraction datasets from Ant Group, to compile a comprehensive training dataset consisting of 2.5 million instances. Subsequently, we undertake full-parameter fine-tuning of the alpaca2-chinese-13b [^5] https://huggingface.co/hfl/chinese-alpaca-2-13b. model on this training dataset, resulting in the refined model named OneKE.
@@ C.2 OneKE
Zero-shot Dataset
To ensure the validity of the zero-shot evaluation and prevent result bias due to data similarity, we select datasets primarily derived from news and biomedical fields as our training sets. This selection is intended to train the model’s capability for instruction following and schema-based extraction. For the evaluation data, we adopt the 13 cross-domain datasets recommended in IEINSTRUCTIONS and YAYI-UIE, which include: for Named Entity Recognition (NER) tasks, we use the CrossNER  (Liu et al., 2021), Weibo NER  (Peng and Dredze, 2015), and Boson [^6] https://github.com/InsaneLife/; in Relation Extraction (RE) tasks, we choose FewRel  (Han et al., 2018), Wiki-ZSL  (Chen and Li, 2021), COAE2016777https://github.com/Sewens/COAE2016, IPRE  (Wang et al., 2019), and SKE2020888https://aistudio.baidu.com/datasetdetail/177191; and for Event Extraction (EE), we include RAMS  (Ebner et al., 2020), WikiEvents  (Li et al., 2021), CrudeOilNews  (Lee et al., 2022b), FewFC  (Zhou et al., 2021), and CCF law  [^9] https://aistudio.baidu.com/projectdetail/4201483. These datasets cover a wide range of fields including literature, music, law, and oil news. It is noteworthy that these evaluation data sets are not used during the training, ensuring that our evaluation accurately reflects the model’s generalization and adaptation capabilities for unseen domains and unseen schema data in zero-shot information extraction.
C.3 Zero-shot performance on Event Extraction
@@
As illustrated in Table  3, the model trained with IEPile exhibits outstanding performance in zero-shot event extraction (EE) tasks, surpassing other baselines. Notably, in the Chinese EE task, the LLaMA2-IEPile model’s performance is slightly inferior to YAYI-UIE’s, revealing LLaMA2’s limitations in processing Chinese data. However, in the English EE task, LLaMA2-IEPile’s performance is significantly superior to that of similar models. This contrast highlights the potential influence of language type on model performance.
Hyperparameter Value
Number of Epochs 5
Learning Rate 5e-5
Batch Size 20
Accumulate 4
Optimizer Adam
Lora_r 64
Lora_alpha 64
Lora_dropout 0.05
Table 4: Training Hyperparameters
C.4 Hyper-parameter
@@
In our research, we select four pre-trained models, Baichuan2-13B-Chat and LLaMA2-13B-Chat, Qwen1.5-14B-Chat, and LLaMA3-8B-Instruct, as the base models for our study. Specifically, we employ the LoRA  (Hu et al., 2022) technique and utilize 8 NVIDIA A800 GPUs to perform instruction tuning on our IEPile dataset. Detailed configurations of the hyperparameters during the fine-tuning process are presented in Table 4.
Dataset Supervised Zero-shot
ACE2004 84.28 77.01
People Daily 98.34 95.29
Table 5: The results of individual LoRA fine-tuning on ACE2004 and People Daily datasets for Baichuan2-13B-Chat, compared with the zero-shot generalization results of Baichuan2-IEPile on these two datasets.
C.5 Supervision Results
Due to limited computational resources, I report only the supervised results for the Baichuan2-IEPile, LLaMA2-IEPile, and OneKE models. Tables  6,  7, and  8 present our experimental results under a supervised learning setting on the training dataset. Specifically, it can be observed that after training on the IEPile, the model excels in Named Entity Recognition (NER), Relation Extraction (RE), and Event Detection (ED), ranking top 2 across these tasks. The model’s performance is only slightly behind other baselines in the Event Argument Extraction. Additionally, we record the model’s performance in Chinese NER, RE, and EE tasks, where it demonstrates robust results. In a comprehensive assessment, the IEPile-trained model showcases performance on par with other models in instruction-based information extraction (IE) tasks and significantly improves performance in zero-shot IE tasks compared to other models. This indicates the significant application prospects and potential of IEPile in the current field of IE.
C.6 Impact of Potential Dataset Bias on Model Performance and Generalization
During the research, we identify that potential biases introduced by the datasets used can affect the model’s performance and generalization capability. Firstly, biases in the definition of schemas within the datasets have a negative impact on model performance  (Huang et al., 2024). In the early stages of training, we observe instability in results due to mutual interference among multiple datasets that contain the same schemas but with differing definitions. For instance, despite wikiann, wikineural, polyglot-NER, and CoNLL2003 all containing common schemas such as people and organization, they each possess distinct scheme definitions. Consequently, in the later stages, only CoNLL2003 is retained. Secondly, the model demonstrates good generalization when dealing with datasets having schemas similar to those in the training set. As shown in Table 5, despite not being included in the training corpus, the People Daily and ACE2004 NER datasets share similar schemas with the MASR and ACE2005 NER dataset in the training set, and the Baichuan2-IEPile model is still capable of handling them proficiently. Lastly, the use of common, coarse-grained labels (such as “person” and “organization”) within the IEPile lead the model, after training, to favor these coarse categories over fine-grained ones (such as “scientist” and “company”) when predicting instructions that included both levels of granularity.
Dataset InstructUIE YAYI-UIE Baichuan2-IEPile LLaMA2-IEPile OneKE
ACE2005 86.66 81.78 81.86 81.14 83.45
AnatEM 90.89 76.54 87.21 86.90 87.88
BC2GM 85.16 82.05 80.73 83.07 82.05
BC4CHEMD 90.30 88.46 90.45 90.07 90.56
BC5CDR 89.59 83.67 88.07 88.01 88.45
CoNLL2003 92.94 96.77 92.49 92.98 93.04
FabNER 76.20 72.63 77.07 76.33 81.06
FindVehicle 89.47 98.47 98.49 97.91 99.45
GENIA-Ent 74.71 75.21 76.66 77.32 78.29
HarveyNER 88.79 69.57 67.70 62.64 69.87
MIT Movie 89.01 70.14 88.23 89.54 89.96
MIT Restaurant 82.55 79.38 79.85 81.30 79.89
MultiNERD 92.32 88.42 94.60 94.24 94.69
NCBI-Disease 90.23 87.29 85.26 87.59 86.95
Ontonotes 90.19 87.04 87.55 90.34 89.08
Avg 87.27 82.49 85.08 85.29 86.24
MSRA - 95.57 87.99 86.32 89.02
Resume NER - - 93.92 92.86 95.84
CLUE NER - - 80.19 76.57 78.43
@@ C.6 Impact of Potential Dataset Bias on Model Performance and Generalization
Dataset InstructUIE YAYI-UIE Baichuan2-IEPile LLaMA2-IEPile OneKE
ADE Corpus 82.31 84.14 83.73 85.87 87.24
CoNLL2004 78.48 79.73 72.87 73.71 76.16
GIDS 81.98 72.36 74.71 74.13 76.69
KBP37 36.14 59.35 65.09 61.49 65.23
NYT 90.47 89.97 93.00 92.22 94.04
NYT11-HRL 56.06 57.53 53.19 54.86 55.56
SciERC 45.15 40.94 43.53 44.58 45.89
Semeval-RE 73.23 61.02 58.47 57.61 61.46
Avg 67.98 68.13 68.07 68.06 70.28
CMeIE - - 49.16 47.40 49.54
DuIE2.0 - 81.19 75.61 74.34 75.73
@@
Dataset InstructUIE YAYI-UIE Baichuan2-IEPile LLaMA-IEPile OneKE
Trigger ACE2005 77.13 65.00 72.46 70.63 71.17
CASIE 67.80 63.00 60.07 61.27 63.82
PHEE 70.14 63.00 66.22 68.52 68.60
Avg 71.69 63.67 66.25 66.81 67.86
DuEE1.0 - 85.00 86.73 84.01 85.75
DuEE-fin - 82.50 83.54 79.00 82.91
Argument ACE2005 72.94 62.71 63.90 62.69 62.75
CASIE 63.53 64.23 56.07 56.78 57.16
PHEE 62.91 77.19 70.85 71.33 72.84
Avg 66.46 68.04 63.60 63.61 64.25
DuEE1.0 - 79.08 75.63 73.79 75.40
DuEE-fin - 70.02 79.34 73.08 78.98
@@
Table 9: Statistical data of Named Entity Recognition (NER) datasets, with an $\dagger$ indicating the zero-shot evaluation set not included in the training. CrossNER  (Liu et al., 2021) is divided into five subsets for our statistical analysis.
@@
COAE2016$\dagger$ 7 General 9 - - 971
@@
SKE2020$\dagger$ 8 News 49 - - 3601
Table 10: Statistical data of Relation Extraction (RE) datasets, with an $\dagger$ indicating the zero-shot evaluation set not included in the training. The test sets for CMeIE and DuIE2.0 are not open-sourced, thus we use the validation sets as our evaluation set. For the FewRel and Wiki-ZSL datasets, we follow  Chia et al. (2022).
@@
CCF law $\dagger$9 Law 9(39) - - 971
Table 11: Statistical data of Event Extraction (EE) datasets, with an $\dagger$ indicating the zero-shot evaluation set not included in the training. The test sets for DuEE1.0 and DuEE-Fin are not open-sourced, thus we use the validation sets as our evaluation set.
Task Instruction & Output
NER 1{ 2 "instruction": "You are an expert in named entity recognition. Please extract entities that match the schema definition from the input. Return an empty list if the entity type does not exist. Please respond in the format of a JSON string.", 3 "schema": ["location", "else", "organization", "person"], 4 "input": "The objective of the Basic Course on War is to provide for combatants of the EPR basic military knowledge for the armed conflict against the police and military apparatus of the bourgeoisie." 5} 6output = { 7 "location": [], 8 "else": [], 9 "organization": ["EPR"], 10 "person": [] 11}
RE 1{ 2 "instruction": "You are an expert in relationship extraction. Please extract relationship triples that match the schema definition from the input. Return an empty list for relationships that do not exist. Please respond in the format of a JSON string.", 3 "schema": ["place of birth", "country capital", "country of administrative divisions", "company"], 4 "input": "Born on May 1 , 1927 , in Brichevo , Bessarabia in the present-day Republic of Moldova , Mr. Bertini emigrated to Palestine with his family as a child and pursued musical studies there , in Milan , and in Paris , where he worked with Nadia Boulanger and Arthur Honegger." 5} 6output = { 7 "place of birth": [{"head": "Mr. Bertini", "tail": "Paris"}], 8 "country capital": [], 9 "country of administrative divisions": [], 10 "company": [] 11}
EE 1{ 2 "instruction": "You are an expert in event extraction. Please extract events from the input that conform to the schema definition. Return an empty list for events that do not exist, and return NAN for arguments that do not exist. If an argument has multiple values, please return a list. Respond in the format of a JSON string.", 3 "schema": [{"event_type": "pardon", "trigger": true, "arguments": ["defendant"]}, {"event_type": "extradite", "trigger": true, "arguments": ["person", "agent", "destination", "origin"]}, {"event_type": "sue", "trigger": true, "arguments": ["place", "plaintiff"]}, {"event_type": "start position", "trigger": true, "arguments": ["person", "entity", "place"]}], 4 "input": "Ethical and legal issues in hiring Marinello" 5} 6output = { 7 "pardon": [], 8 "extradite": [], 9 "sue": [], 10 "start position": [{"trigger": "hiring", "arguments": {"person": "Marinello", "entity": "NAN", "place": "NAN"}}] 11}